OpenStudy (anonymous):

how do i simplify 5 square root 320

OpenStudy (anonymous):

\[5* \sqrt{320} -- or -- \sqrt[5]{320}\]

OpenStudy (anonymous):

5√320

OpenStudy (anonymous):

5 \sqrt{320} = 5 \sqrt{2*160} = 5 \sqrt{2*2*80} = 10 \sqrt{80} = 10 \sqrt{2*2*20} = 20 \sqrt{20} = 20 \sqrt{2*2*5} = 40 \sqrt{5}

OpenStudy (anonymous):

5√{320} = 5√{2*2*80} = 10√{80} = 20√{20} = 40√{5}

OpenStudy (anonymous):

Does this make sense?